# Venkata Ramana Makkapati Honda Aircraft Company

### v1

@@ Abstract
This paper introduces a novel reachability problem for the scenario where one agent follows another agent using the constant bearing pursuit strategy, and analyzes the geometry of the reachable set of the follower. Key theoretical results are derived, providing bounds for the associated dependent reachable set. Simulation results are presented to empirically establish the shape of the dependent reachable set. In the process, an original optimization problem for the constant bearing strategy is formulated and analyzed.
@@ 1 Introduction
Reachability theory is widely applied to address collision avoidance and safety in the planning and control of multi-agent systems [althoff2014online, malone2017hybrid, wang2020infusing]. The reachability analysis involves identifying the set that contains all the states an agent can reach given the initial conditions and the allowable control inputs, as a function of time. In instances where one agent follows another agent using a feedback strategy, our problem of interest is to identify the follower’s reachable set, which is referred to as the dependent reachable set (DRS) in this paper. The idea of a DRS was introduced by Comandur et al. to analyze deception relevance in games with asymmetric information [comandur2024desensitization]. The problem becomes relevant when the leader (the agent being followed) is a faulty robot [park2017fault], or when the leader itself is the opponent in strategic engagements [ho2022game].
The reachability problem, a dual of the controllability problem, was extensively studied for linear systems [gayek1991survey]. Computing reachable sets for nonlinear continuous-time systems is more complex, as an analytical expression for the set is not always guaranteed [althoff2021set]. Consequently, several approaches including those based on optimal control theory [mitchell2005time, chen2018hamilton], barrier functions [kong2018reachable], and trajectory sensitivity analysis [ramdani2008reachability, chuchu2016locally], were developed to estimate reachable sets. More recently, there has been an increased focus on reachability analysis for multi-agent systems [thapliyal2024algorithm, wang2021formal], and neural network controllers [gates2023scalable, everett2021reach].
The constant bearing pursuit strategy is also known as “parallel navigation” in the robotics literature [strydom2015biologically], and is popular among missile guidance problems [Shneydor], and differential games [isaacs1999differential]. The guidance strategy was applied to mobile robots to reach a moving target [belkhouche2007parallel, rafie2010time].
This paper characterizes DRS for an agent that employs the constant bearing strategy to pursue another agent. To the best of authors’ knowledge, the concept of DRS has not been studied, and is new to the literature. The contributions of this paper are as follows: a) the concept of DRS is introduced for multi-agent systems; b) the geometry of DRS is characterized for an agent that employs the constant bearing pursuit strategy; c) a novel optimization problem for scenarios involving the constant bearing strategy is formulated and discussed.
@@
The paper is organized as follows. Section 2 presents the problem formulation and the underlying assumptions. Section 3 discusses the theory of reachable sets. Section 4 provides key results, both theoretical and simulation, that describe the geometry of DRS for the constant bearing strategy. Section 5 studies the optimization problem to find the closest and farthest points from the leader that the agent can reach while employing the constant bearing strategy. Section 6 concludes the paper and provides some directions for future work.
@@ 2 Problem Statement
Consider a scenario involving two agents (denoted using the subscripts $I$ and $D$) moving in the Cartesian plane at a constant speed by controlling their respective heading angles. The agents’ dynamics can be expressed as
@@
Without loss of generality, we consider $\mathbf{x}^{0}_{D}$ to be the origin, and $\mathbf{x}^{0}_{I}=[a,0]^{\top}$. Consequently, the initial line of sight (LOS) is parallel to the horizontal axis. Since the dependent agent follows the constant bearing pursuit strategy, it chooses a heading angle such that the LOS does not rotate, and the relative distance from the independent agent is reduced at a given time instant [makkapati2018optimal]. Therefore,
@@
In the remainder of this paper, we consider only the case where $\mathbf{u}_{D}(t)=[v_{D}\cos\psi_{D}(t),~v_{D}\sin\psi_{D}(t)]^{\top}$ for $\psi_{D}(t)$ given in (2)
@@
Given the dynamics of the independent agent per (1), its trajectory $\mathbf{x}_{I}(t),~t\geq 0$ is a function of the control function $\mathbf{u}_{I}\in\mathcal{U}_{I}$. For every trajectory of the independent agent, there is a corresponding trajectory of the dependent agent that is obtained using the constant bearing feedback strategy. Now, the problem statement examined in this paper is presented below.
@@
For the set of all feasible trajectories of the independent agent, determine the corresponding set, $\mathcal{D}(t)$, of positions that will be reached by the dependent agent.
In other words, given that the dependent agent is committed to a constant bearing pursuit strategy, what is the set of points that the independent agent can force the dependent agent to reach? When $v_{D}>v_{I}$, every possible trajectory of the independent agent will result in a capture in finite time by the dependent agent that follows the constant bearing pursuit strategy. Here, capture refers to the event $\|\mathbf{x}_{D}(t)-\mathbf{x}_{I}(t)\|_{2}\approx 0$. In this regard, the paper considers that the set $\mathcal{D}(t)$ comprises only those points that correspond to the active pursuit trajectories at a given time. By active pursuit trajectories, we mean those trajectories of the dependent agent that are yet to result in capture for the corresponding trajectory of the independent agent. The maximum capture time $t_{c}=a/(v_{D}-v_{I})$ corresponds to an instance when the independent agent chooses the control input $\mathbf{u}_{I}(t)=[v_{I},0]^{\top}$ for all time until capture. In the pursuit-evasion literature, this chosen control input of the independent agent corresponds to what is known as pure evasion, which guarantees the maximum capture time. Note that as $v_{I}\rightarrow v_{D}$, $t_{c}\rightarrow\infty$.
@@ 3 Reachable Sets
A standard definition for the independent agent’s reachable set at time $t\geq 0$ is given below.
@@
The independent agent’s reachable set, $\mathcal{R}_{I}(\mathbf{x}_{I}^{0},t)$, at time $t\geq 0$ with the initial state at $\mathbf{x}_{I}^{0}$ is the set of all points that the agent can reach at time $t$:
@@
A closed-form expression for the reachable set of the independent agent, per the dynamics in (1), can be obtained using the following lemma. Hereafter, the dependency of the reachable set on the initial condition, which is apparent, is dropped for brevity, and the reachable set at time $t$ is denoted as $\mathcal{R}_{I}(t)$.
@@
Without loss of generality, let $\mathbf{x}_{I}^{0}$ be the origin. Using Lemma 3.2, all points on the line segment joining the origin and the point $v_{I}T\cos\theta,v_{I}T\sin\theta]^{\top}$ can be reached with the control function of the form
@@ 4 Dependent Reachable Sets
The solution to Problem 2.2 involves identifying a specialized reachable set of the dependent agent, termed the dependent reachable set (DRS). DRS is defined for the agent that determines its control input using a feedback control function of the form $\mathbf{u}_{D}(t)=f(\mathbf{x}_{D}(t),\mathbf{x}_{I}(t),\mathbf{u}_{I}(t))$, which is a function of the instantaneous state and control input of the independent agent. A formal definition for the DRS is presented below.
@@
$\mathbf{x}=\mathbf{x}_{D}^{0}+\int_{0}^{t}f(\hat{\mathbf{x}}_{D}(\tau),\hat{\mathbf{x}}_{I}(\tau),\hat{\mathbf{u}}_{I}(\tau))\mathrm{d}\tau\Bigg\},$ (9)
@@
In this paper, we analyze dependent reachable sets in the case where the dependent agent follows the constant bearing strategy. We denote the DRS by $\mathcal{D}(t)$, dropping the dependencies on the initial conditions $\mathbf{x}_{D}^{0}$, $\mathbf{x}_{I}^{0}$ for the sake of notation brevity.
For the problem described in Section 2, the reachable set of the independent agent $\mathcal{R}_{I}(t)$ is a circular region with its center at $\left(a,0\right)$ and radius $v_{I}t$ (see Lemma 3.3). Similarly, the reachable set of the dependent agent $\mathcal{R}_{D}(t)$ is also a circle with its center at the origin and radius $v_{D}t$:
@@
$\mathcal{D}(t)\subseteq\mathcal{R}_{D}(t).$ (11)
Figure 1: Schematics of the two distinct scenarios that are considered to characterize the dependent reachable set for the constant bearing strategy
@@ 4.1 Theoretical Results
This subsection presents the characterization of the DRS for the constant bearing strategy at time $t$, including the points $P_{1}\left(t\sqrt{v_{D}^{2}-v_{I}^{2}},v_{I}t\right)$ and $P_{2}\left(t\sqrt{v_{D}^{2}-v_{I}^{2}},-v_{I}t\right)$ lying on $\partial\mathcal{R}_{D}(t)$, which are obtained as a result of the constant bearing strategy attributes.
As shown in Fig. 1, two different scenarios are considered to develop proofs that provide the shape of the DRS at a given instant in time $t\in[0,t_{c}]$. The vertical diameter of $\mathcal{R}_{I}(t)$ (depicted using a dashed line in Fig. 1) is the line joining $(a,v_{I}t)$ and $(a,-v_{I}t)$, which is perpendicular to the initial LOS (the horizontal axis). The first scenario in Fig. 1 depicts the case where the vertical diameter of $\mathcal{R}_{I}(t)$ does not completely lie in the interior of the circle $\mathcal{R}_{D}(t)$. The first scenario could correspond to the circle $\mathcal{R}_{D}(t)$ either intersecting or not intersecting the vertical diameter of $\mathcal{R}_{I}(t)$. The earliest time instant when $\partial\mathcal{R}_{D}(t)$ first intersects the vertical diameter is at $t_{1}=a/v_{D}$. Note that at time $t_{1}$, the initial position of the independent agent is part of the boundary of the dependent agent’s reachable set. At the same time, it can be deduced that the latest time instant when $\partial\mathcal{R}_{D}(t)$ intersects the vertical diameter is at $t_{2}=a/\sqrt{v_{D}^{2}-v_{I}^{2}}$. In summary, the first scenario in Fig. 1 corresponds to instances occurring in the time interval $0\leq t<t_{2}$.
Consequently, the second scenario in Fig. 1 depicts the case where $\partial\mathcal{R}_{D}(t)$ fully encompasses the vertical diameter of $\mathcal{R}_{I}(t)$. The instances corresponding to the second scenario occur in the time interval $t_{2}<t\leq t_{c}$. The following lemmas help characterize the DRS for the constant bearing strategy.
Figure 2: Characterization of the DRS for $0\leq t\leq t_{2}$
@@
For the constant bearing strategy, the LOS, which is initially aligned along the horizontal axis, does not rotate until capture. As a result, $y_{D}(t)=y_{I}(t)$ for all time $t\in[0,t_{c}]$, and the vertical coordinate of the dependent agent is bounded by $\pm v_{I}t$. Therefore,
@@
Given $\psi_{D}(t)$ is the heading angle of the dependent agent at time $t$, per (2),
@@
Consequently, at time $t$, the minimum horizontal speed of the dependent agent is
$\underset{\psi_{I}(t)\in(-\pi,\pi]}{\min}~v_{D}\cos\psi_{D}(t)=\sqrt{v_{D}^{2}-v_{I}^{2}}.$ (14)
As a result, the minimum horizontal distance that the dependent agent can travel within time $t$ is $t\sqrt{v_{D}^{2}-v_{I}^{2}}$. Note that the dependent agent always has a non-negative horizontal speed since $\psi_{D}(t)\in\left[-\dfrac{\pi}{2},\dfrac{\pi}{2}\right]$, for $t\in[0,t_{c}]$ and $v_{D}\geq v_{I}$. Therefore,
@@
From Lemmas 4.2 and 4.3, it can be deduced that $\mathcal{D}(t)$ will lie in the intersection of the regions given in (10), (12), and (15). The boundaries of these three regions intersect at the points $P_{1}$ and $P_{2}$. The intersection region for the first scenario ($0\leq t\leq t_{2}$) is shown in Fig. 2. Note that at time $t_{2}$, the line joining the points $P_{1}$ and $P_{2}$ (which lie on $\partial\mathcal{R}_{D}(t)$) coincides with the vertical diameter of $\mathcal{R}_{I}(t)$. The following theorem establishes that the DRS for the first scenario is the intersection region itself.
@@
Expressions in (10) and (15) establish that the sets $\|\mathbf{x}\|_{2}\leq v_{D}t$ and $x\geq t\sqrt{v_{D}^{2}-v_{I}^{2}}$ bound the DRS for all time $t\in[0,t_{c}]$. The proof involves showing that for every point $\mathbf{x}$ from the intersection of the aforementioned sets, there exists a control function $\mathbf{u}_{D}(.)\in\mathcal{U}_{D}$ which drives the dependent agent following the constant bearing strategy to that $\mathbf{x}$. In this regard, for time $t\in(0,t_{2}]$ and $\theta\in\left[-\dfrac{\pi}{2},\dfrac{\pi}{2}\right]$, consider the switching control function of the type
$\mathbf{u}_{2s}(\tau)=\begin{cases}[v_{I}\cos\theta,v_{I}\sin\theta]^{\top},\quad&\text{if }0\leq\tau<t_{s},\\ \dfrac{\pi}{2},\quad&\text{if }t_{s}\leq\tau<\dfrac{t+t_{s}}{2},\\ -\dfrac{\pi}{2},\quad&\text{if }\dfrac{t+t_{s}}{2}\leq\tau\leq t,\end{cases},$ (17)
for the independent agent. Figure 3 indicates the resulting path (lines with arrows depicting the direction of motion) of the independent agent for the control function in (17). The corresponding horizontal coordinate of the dependent agent can be obtained as
@@
The corresponding vertical coordinate for both agents can be obtained as $t_{s}v_{I}\sin\theta$.
For a constant vertical coordinate $y=t_{s}v_{I}\sin\theta$, $t_{s}\in[y/v_{I},t]$, the dependent agent’s horizontal coordinate in (18) can be rewritten as
@@
From (19), it can be observed that all the points from the intersection region in (16) for a given $y\in[-v_{I}t,v_{I}t]$ are spanned by $t_{s}\in[y/v_{I},t]$, and correspond to the points reached by the independent agent using the control function in (17) for $\pi/2\leq\theta\leq\theta_{l}$ (when $y\geq 0$) or $\theta_{l}\leq\theta\leq-\pi/2$ (when $y<0$). Note that $\theta_{l}=\tan^{-1}(y/v_{I}t)$
@@
Next, we show that all the resulting trajectories of the dependent agent corresponding to the control function in (17), and $\theta\in\left[-\dfrac{\pi}{2},\dfrac{\pi}{2}\right]$ will not result in capture for $0\leq t<t_{2}$. For a capture to occur at time $t$ with the independent agent’s control function in (17), the horizontal coordinates of both agents should be the same. Consequently,

### v2

# Dependent Reachable Sets for the Constant Bearing Pursuit Strategy

@@ Abstract
This paper introduces a novel reachability problem for the scenario involving two agents, where one agent follows another agent using a feedback strategy. The geometry of the reachable set for an agent, termed dependent reachable set, is characterized using the constant bearing pursuit strategy as a case study. Key theoretical results are presented that provide geometric bounds for the associated dependent reachable set. Simulation results are presented to empirically establish the shape of the dependent reachable set. In the process, an original optimization problem is formulated and analyzed for the constant bearing pursuit strategy.
Keywords: Reachability, Feedback Strategy, Multi-agent Systems, Apollonius Circle, Constant Bearing Strategy, Switching Control
@@ 1 Introduction
Reachability analysis is a key formal verification tool for autonomous and cyber-physical systems [1, 9, 12]. It provides provable safety guarantees and enables fault-tolerant control under uncertainties [5, 11]. Forward reachability analysis computes the forward reachable set, which contains all states a system (or agent) can reach over time, given the initial state and the set of allowable control inputs. In contrast, backward reachability analysis identifies the backward reachable set, the set of states from which a given target state can be reached at a specified time. In general, it is challenging to analytically compute the forward or backward reachable sets for nonlinear systems, even for most constrained linear systems [2]. Consequently, several approaches, including those based on optimal control theory [17, 4, 23], barrier functions [13], zonotopes [14, 26], and trajectory sensitivity analysis [20, 8], were developed to numerically estimate reachable sets for linear and nonlinear systems. In recent years, there has been an increased focus on reachability analysis for multi-agent systems [24, 25] and neural network controllers [10, 7].
In the context of formal verification for multi-agent systems, where one agent (referred to as dependent agent) follows another agent (referred to as independent agent) using a feedback strategy, it is important to characterize the reachable set of the dependent agent for a given strategy, termed dependent reachable set (DRS) in this paper. This problem is particularly relevant in defense and security applications, where it is critical to estimate the worst-case scenario for the dependent agent by evaluating the set of states that the agent can reach when following an adversary using a feedback strategy. Conversely, to plan for contingencies in strategic engagements, the independent agent must identify the set of locations to which it can lead the dependent agent. To this end, in this paper, DRS for two-dimensional problems involving two agents is formalized, which can be easily extended to systems with more than two agents and to three-dimensional environments.
Regarding the applicability of DRS, Comandur et al. noted that for a class of pursuit-evasion games involving deception, estimating such reachable sets for the deceptive player (which in our paper corresponds to the independent agent) is essential to assess the relevance of its strategy [6]. Their work considered the pure pursuit strategy for the dependent agent. However, the lack of closed-form solutions in the existing literature on pure pursuit strategy [21, 15] and our preliminary analysis suggest that the geometry of the associated DRS cannot be expressed analytically in closed form. We therefore focus on the constant bearing pursuit strategy, which is widely used in missile guidance problems [21]. The constant bearing pursuit strategy is also known as “parallel navigation” in the robotics literature [22], and was applied to mobile robots to reach a moving target [3, 18, 16]. The geometric elegance of constant bearing pursuit is supported by its connection to the Apollonius circle [15], a property also evident in our results.
A basic approach to computing the DRS at a fixed time $t$ is to first determine the set of states reachable by the dependent agent for each possible state of the independent agent at the same time $t$, and then perform a union of these sets. Note that the independent agent may reach a certain state via multiple trajectories, allowing the dependent agent to reach multiple corresponding states. In this regard, an optimization problem is formulated to identify the maximum and minimum relative distances that the dependent agent can achieve, given the initial conditions and current position of the independent agent, at a specified time, while following the constant bearing pursuit strategy. Although an analytical solution to this optimization remains elusive, simulations reveal interesting geometric properties linked to ellipses.
To the best of the authors’ knowledge, the concept of DRS is novel and unexplored in the multi-agent systems literature. The contributions of this paper are as follows.
  1. 1.
    The concept of DRS is formalized.
  2. 2.
    The geometry of DRS for the constant bearing pursuit strategy is characterized using theoretical and simulation results.
  3. 3.
    A theoretical result that establishes the link between the DRS of the constant bearing pursuit strategy and the Apollonius circle is presented.
  4. 4.
    An original optimization problem for maximum and minimum relative distances under the constant bearing pursuit strategy is formulated and analyzed.
  5. 5.
    A new property of ellipses is observed in the context of the constant bearing pursuit strategy using empirical results.
@@
The remainder of the paper is organized as follows. Section 2 presents the problem formulation and the underlying assumptions. Section 3 reviews the theory of reachable sets. Section 4 presents key theoretical and simulation results describing the geometry of the DRS for constant bearing pursuit. Section 5 analyzes the optimization problem to find the closest and the farthest points from the independent agent that the dependent agent can reach while employing the constant bearing pursuit strategy. Section 6 concludes the paper and outlines directions for future work.
@@ 2 Problem Statement
Consider the standard two-dimensional scenario involving two agents (denoted using the subscripts $I$ and $D$) moving in the Cartesian plane at a constant speed by controlling their respective heading angles. The agents’ dynamics can be expressed as
@@
Without loss of generality, we set $\mathbf{x}^{0}_{D}=[0,0]^{\top}$ (origin) and $\mathbf{x}^{0}_{I}=[a,0]^{\top}$. Consequently, the initial line of sight (LOS) is parallel to the horizontal axis. Since the dependent agent follows the constant bearing pursuit strategy, it chooses a heading angle at a given time instant such that the LOS does not rotate, i.e., the LOS is parallel to the horizontal axis for all time, and the relative distance between the agents does reduce [15]. Therefore,
@@
In the remainder of this paper, we consider only the case where $\mathbf{u}_{D}(t)=[v_{D}\cos\psi_{D}(t),~v_{D}\sin\psi_{D}(t)]^{\top}$ for $\psi_{D}(t)$ given in (2).
@@
Given the dynamics of the independent agent per (1), its trajectory $\mathbf{x}_{I}(t),~t\geq 0$ is a function of the control input $\mathbf{u}_{I}\in\mathcal{U}_{I}$. Since $\mathbf{u}_{D}(t)$ is a function of $\mathbf{u}_{I}(t)$, per (2), for every trajectory of the independent agent, there is a corresponding trajectory of the dependent agent. Now, the problem statement examined in this paper is presented below.
@@
For the set of all feasible trajectories of the independent agent, determine the corresponding set, $\mathcal{D}(t)$, of positions (or states) that will be reached by the dependent agent following the constant bearing pursuit strategy.
In other words, given that the dependent agent is committed to the constant bearing pursuit strategy, what is the set of points to which the independent agent can drive the dependent agent? When $v_{D}>v_{I}$, every possible trajectory of the independent agent will result in a capture in finite time by the dependent agent that follows the constant bearing pursuit strategy. Here, capture refers to the event $\|\mathbf{x}_{D}(t)-\mathbf{x}_{I}(t)\|_{2}\approx 0$. In this regard, the paper considers that the set $\mathcal{D}(t)$ comprises only those points that correspond to the active pursuit trajectories at a given time. By active pursuit trajectories, we mean those trajectories of the dependent agent that have yet to result in capture for the corresponding trajectory of the independent agent. The maximum capture time $t_{c}=a/(v_{D}-v_{I})$ corresponds to an instance in which the independent agent chooses the control input $\mathbf{u}_{I}(t)=[v_{I},0]^{\top}$ for all time until capture. In the pursuit-evasion literature, this chosen control input of the independent agent corresponds to what is known as pure evasion, which guarantees the maximum capture time. Note that as $v_{D}\rightarrow v_{I}$, $t_{c}\rightarrow\infty$.
@@ 3 Reachable Sets
The standard definition for the reachable set of the independent agent at time $t\geq 0$ is given below.
@@
The reachable set of the independent agent, $\mathcal{R}_{I}(\mathbf{x}_{I}^{0},t)$, at time $t\geq 0$ with the initial state at $\mathbf{x}_{I}^{0}$ is the set of all points that the agent can reach at time $t$:
@@
Hereafter, the dependence of the reachable set on the initial condition, which is apparent, is dropped for brevity, and the reachable set at time $t$ is denoted as $\mathcal{R}_{I}(t)$. A closed-form expression for the reachable set of the independent agent, per the dynamics in (1), can be obtained using the following lemmas. Though the analytical solution for the reachable set is common knowledge, to the best of the authors’ knowledge, a formal proof for the same is not found in the existing literature. In addition, the theoretical results for the dependent reachable set, presented in Section 4, are developed based on the following lemmas.
@@
Without loss of generality, let $\mathbf{x}_{I}^{0}$ be the origin. Based on Lemma 3.2, all points on the line segment that join the origin and the point $v_{I}T\cos\theta,v_{I}T\sin\theta]^{\top}$ can be reached with the control function of the form
@@ 4 Dependent Reachable Sets
The solution to Problem 2.2 involves identifying a specialized reachable set of the dependent agent, termed the dependent reachable set. DRS is defined for the agent that determines its control input using a feedback control function of the form $\mathbf{u}_{D}(t)=f(\mathbf{x}_{D}(t),\mathbf{x}_{I}(t),\mathbf{u}_{I}(t))$, which is a function of the instantaneous state and the control input of the independent agent. A formal definition for the DRS is presented below.
@@
For the problem described in Section 2, the reachable set of the independent agent $\mathcal{R}_{I}(t)$ is a circular region with its center at $\left(a,0\right)$ and radius $v_{I}t$ (see Lemma 3.3). Similarly, the reachable set of the dependent agent $\mathcal{R}_{D}(t)$ is also a circular region with its center at the origin and radius $v_{D}t$:
@@
$\mathcal{D}_{f}(\mathbf{x}_{D}^{0},\mathbf{x}_{I}^{0},t)\subseteq\mathcal{R}_{D}(t).$ (11)
In this paper, dependent reachable sets are analyzed in the situation where the dependent agent follows the constant bearing pursuit strategy. The DRS for the constant bearing pursuit strategy is denoted by $\mathcal{D}(t)$, dropping the dependencies on the initial conditions $\mathbf{x}_{D}^{0}$ and $\mathbf{x}_{I}^{0}$ for brevity.
@@ 4.1 Theoretical Results
This subsection presents theoretical results for the characterization of DRS $\mathcal{D}(t)$. As shown in Figure 1, two different scenarios are considered to develop proofs that characterize the shape of the DRS at a given instant in time $t\in[0,t_{c}]$. In this regard, the line segment joining the points $(a,v_{I}t)$ and $(a,-v_{I}t)$ is denoted as $\mathcal{V}$, which is the set of all points that form the vertical diameter of the circle $\partial\mathcal{R}_{I}(t)$. The line segment $\mathcal{V}$ is perpendicular to the initial LOS (the horizontal axis) and is represented by a dashed line in Figure 1.
The first scenario in Figure 1 depicts the case of $\mathcal{V}\nsubseteq\mathcal{R}_{D}(t)$. In the first scenario, there are two possibilities: 1) the line segment $\mathcal{V}$ is outside the circle $\partial\mathcal{R}_{D}(t)$ (shown in Figure 1); or 2) the line segment $\mathcal{V}$ intersects the circle $\partial\mathcal{R}_{D}(t)$. The first time instant when $\partial\mathcal{R}_{D}(t)$ intersects the line segment $\mathcal{V}$ is at $t_{1}=a/v_{D}$, which is when $\mathcal{V}$ is tangential to the circle $\partial\mathcal{R}_{D}(t)$. Note that at time $t_{1}$, the initial position of the independent agent $a,0]^{\top}$ lies on the circle $\partial\mathcal{R}_{D}(t)$. Subsequently, let $t_{2}$ be the last time instant when the circle $\partial\mathcal{R}_{D}(t)$ intersects the line segment $\mathcal{V}$. Due to the symmetry associated with the problem, which can be observed in Figure 1, it can be deduced that at time $t_{2}$, $\partial\mathcal{R}_{D}(t)$ contains the end points of the line segment $\mathcal{V}$, $(a,v_{I}t)$ and $(a,-v_{I}t)$. As a result, we have
$v_{D}^{2}t_{2}^{2}$ $=a^{2}+v_{I}^{2}t_{2}^{2}$
$\implies t_{2}$ $=\frac{a}{\sqrt{v_{D}^{2}-v_{I}^{2}}}.$ (12)
Therefore, the first scenario, $\mathcal{V}\nsubseteq\mathcal{R}_{D}(t)$, corresponds to instances that occur in the time interval $0\leq t<t_{2}$.
Figure 1: Schematics of the two distinct scenarios that are considered to characterize the dependent reachable set for the constant bearing pursuit strategy
The second scenario in Figure 1 depicts the case where the circle $\partial\mathcal{R}_{D}(t)$ fully encompasses the line segment $\mathcal{V}$ i.e. $\mathcal{V}\subseteq\mathcal{R}_{D}(t)$. The instances corresponding to the second scenario occur in the time interval $t_{2}\leq t\leq t_{c}$. The following lemmas help characterize the DRS for the constant bearing pursuit strategy.
@@
For the constant bearing pursuit strategy, the LOS, which is initially aligned along the horizontal axis, does not rotate. As a result, $y_{D}(t)=y_{I}(t)$ for all time $t\in[0,t_{c}]$, and the vertical coordinate of the dependent agent is bounded by $\pm v_{I}t$. Therefore,
@@
For this proof, the minimum horizontal speed achievable by the dependent agent with the constant bearing pursuit strategy must be identified. To this end, since the dependent agent’s heading ($\psi_{D}(t)$) is a function of the independent agent’s heading ($\psi_{I}(t)$) per (2), the $\psi_{I}(t)$ corresponding to the minimum horizontal speed can be obtained as
@@
Note that $\psi_{I}(t)=\pm\pi/2$ indicates that the independent agent moves perpendicular to the LOS. Consequently, at time $t$, the minimum horizontal speed of the dependent agent is
$\underset{\psi_{I}(t)\in(-\pi,\pi]}{\min}~v_{D}\cos\psi_{D}(t)$ $=v_{D}\cos\Big(\sin^{-1}\left(\dfrac{v_{I}}{v_{D}}\right)\Big)$
$=\sqrt{v_{D}^{2}-v_{I}^{2}}.$ (15)
As a result, the minimum horizontal distance that the dependent agent can travel within time $t$ is $t\sqrt{v_{D}^{2}-v_{I}^{2}}$. Therefore,
@@
From Lemmas 4.2 and 4.3, it can be deduced that $\mathcal{D}(t)$ will lie in the intersection of the regions given in (10), (13), and (16). The intersection region for the first scenario ($0\leq t\leq t_{2}$) is shown in Figure 2. The boundaries of these three regions intersect at the points $P_{1}$ and $P_{2}$, as shown in Figure 2. The coordinates of the points $P_{1}$ and $P_{2}$ can be obtained as $\left(t\sqrt{v_{D}^{2}-v_{I}^{2}},v_{I}t\right)$ and $\left(t\sqrt{v_{D}^{2}-v_{I}^{2}},-v_{I}t\right)$, respectively. From the aforementioned coordinates, it can be seen that the points $P_{1}$ and $P_{2}$ lie on the circle $\partial\mathcal{R}_{D}(t)$. Note that at time $t_{2}$, the line segment joining the points $P_{1}$ and $P_{2}$ coincides with $\mathcal{V}$, the vertical diameter of the circle $\partial\mathcal{R}_{I}(t)$. In Figure 2, $P_{x}=[v_{D}t,0]^{\top}$. The discussion of the Apollonius circle (shown in Figure 2) and its link to the DRS is included after Theorem 4.4. The following theorem establishes that the DRS for the first scenario is the intersection region itself.
Figure 2: Characterization of the DRS for $0\leq t\leq t_{2}$
@@
The expressions in (10) and (16) establish that the sets $\|\mathbf{x}\|_{2}\leq v_{D}t$ and $x\geq t\sqrt{v_{D}^{2}-v_{I}^{2}}$ bound the DRS for all time $t\in[0,t_{c}]$. The proof involves first showing that for every point $\mathbf{x}$ from the intersection of the aforementioned sets, there exists a control function $\mathbf{u}_{I}(.)\in\mathcal{U}_{I}$ of the independent agent which drives the dependent agent following the constant bearing pursuit strategy to the point $\mathbf{x}$. In this regard, for time $t\in(0,t_{2}]$ and $\theta\in\left[-\dfrac{\pi}{2},\dfrac{\pi}{2}\right]$, consider the candidate switching control function of the type
$\mathbf{u}_{2s}(\tau)=\begin{cases}[v_{I}\cos\theta,v_{I}\sin\theta]^{\top},\quad&\text{if }0\leq\tau<t_{s},\\ [0,v_{I}]^{\top},\quad&\text{if }t_{s}\leq\tau<\dfrac{t+t_{s}}{2},\\ [0,-v_{I}]^{\top},\quad&\text{if }\dfrac{t+t_{s}}{2}\leq\tau\leq t,\end{cases},$ (18)
for the independent agent. Since the independent agent starts at $(a,0)$, for the control function in (18), the independent agent reaches the point $\left(a+t_{s}v_{I}\cos\theta,t_{s}v_{I}\sin\theta\right)$ at time $t_{s}$, then travels vertically up until time $\dfrac{t+t_{s}}{2}$, and finally returns to the point $\left(a+t_{s}v_{I}\cos\theta,t_{s}v_{I}\sin\theta\right)$ at time $t$. Figure 3 indicates the resulting path (lines with arrows that depict the direction of motion) of the independent agent for the control function in (18).
For the control function in (18), $\psi_{D}(t)$ for time $t\in[0,t_{s})$ can be obtained as $\sin^{-1}\left(\dfrac{v_{I}\sin\theta}{v_{D}}\right)$ per (2). As a result, the dependent agent reaches the point $\left(t_{s}\sqrt{v_{D}^{2}-v_{I}^{2}\sin^{2}\theta},t_{s}v_{I}\sin\theta\right)$ at time $t_{s}$. Note that the dependent agent starts from the origin. As discussed in the proof of Lemma 4.2, the horizontal speed of the dependent agent for the time when the independent agent moves perpendicular to the LOS (horizontal axis) is $\sqrt{v_{D}^{2}-v_{I}^{2}}$. Therefore, at time $t$, the dependent agent reaches the horizontal coordinate
@@
Note that the corresponding vertical coordinate is $t_{s}v_{I}\sin\theta$, which is the same as that of the independent agent.
Given $y=t_{s}v_{I}\sin\theta$, the dependent agent’s horizontal coordinate in (19) can be rewritten as
@@
From (20), it can be observed that all the points from the set in (17) are spanned using $t_{s}\in[y/v_{I},t]$, given $y\in[-v_{I}t,v_{I}t]$. Each point from the set in (17) corresponds to the point reached by the independent agent using the control function in (18) for $\pi/2\leq\theta\leq\theta_{l}$ (when $y\geq 0$) or $\theta_{l}\leq\theta\leq-\pi/2$ (when $y<0$), where $\theta_{l}=\sin^{-1}(y/v_{I}t)$.
@@
Next, we need to show that all the resulting trajectories of the dependent agent corresponding to the control function in (18), and $\theta\in\left[-\dfrac{\pi}{2},\dfrac{\pi}{2}\right]$ will not result in capture for $0\leq t<t_{2}$. For a capture to occur at time $t$ with the independent agent’s control function in (18), both agents should reach the same horizontal coordinate at time $t$. In this regard, for the capture at $t$,